OpenStudy (anonymous):

Select the equation of the line that passes through the point (3, -1) and is parallel to the line on the graph. y = -1 y = 3 y = x -1 y = 3x - 1

OpenStudy (anonymous):

@PrettyGirlSwag<3

OpenStudy (anonymous):

@PrettyGirlSwag<3

OpenStudy (anonymous):

If two lines are parallel they will have the same slope. So first you need to find the slope of the given line 2y - 16x = 3. . To find the slope you can work this equation into the slope intercept form of y = mx + b. When you get it into this form you will know that the slope is m ... the constant that multiplies the x. . Start with: . 2y - 16x = +3 . get rid of the -16x on the left side by adding 16x to both sides to get: . 2y = +16x + 3 . Solve for y by dividing all terms on both sides by 2 ... the multiplier of y. When you do this division you get: . y = (16/2)*x + 3/2 . In the first term on the right side the division 16/2 = 8. So the equation becomes: . y = 8x + 3/2 . Note that this is now in the slope intercept form of y = mx + b. Therefore, the multiplier of the x term is the slope. So the slope is +8. Any line that is parallel to this line must have a slope of 8. . So far we know that the slope of a parallel line must be of the form: . y = 8x + b

OpenStudy (anonymous):

soooooooo what is the answer if the slope is y=8x+b

OpenStudy (anonymous):

wish I could help you >.>

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what is the answer tho